OpenStudy (anonymous):

find an equation having -2, i, 3i as solutions

OpenStudy (anonymous):

a polynomial equation?

OpenStudy (anonymous):

(x+2)(x-i)(x-3i)=0 A polynomial equation in x

OpenStudy (anonymous):

would look like \[(x+2)(x-i)(x-3i)\] but if you want real coefficients then you need the conjugates as well \[(x+2)(x-i)(x+i)(x-3i)(x+3i)\] will do it

OpenStudy (anonymous):

\((x-i)(x+i)=x^2+1\) and \((x-3i)(x+3i)=x^2+9\) so if you want real coefficients multiply out \[(x+2)(x^2+1)(x^2+9)\]

OpenStudy (anonymous):

thanks guys for your help

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

too bad i could only choose one for the best answer but appreciate satellite73 answers very much for explaining everything that helped my daughter enhanced her knowledge more about it. thank you once again satellite73 and su225